OpenStudy (anonymous):

show that matrix | 1 a b ab | | 1 a' b a'b | | 1 a b' ab' | | 1 a' b' a'b | + = (a'-a)(b'-b)^2

OpenStudy (badhi):

are you sure the (4,4) element is correct shouldn't it be a'b'

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the problem ?

OpenStudy (badhi):

$$\left|\begin{matrix} 1 & a & b & ab\\ 1 & a' & b & a'b\\ 1 & a & b' & ab'\\ 1 & a' & b' & a'b \end{matrix}\right|\xrightarrow[R_2-R_4\rightarrow R_2]{R_1-R_3\rightarrow R_1}$$ $$\left|\begin{matrix} 0 & 0 & (b-b') & a(b-b')\\ 0 & 0 & (b-b') & 0\\ 1 & a & b' & ab'\\ 1 & a' & b' & a'b \end{matrix}\right|=(b-b')^2\left|\begin{matrix} 0 & 0 & 1 & a\\ 0 & 0 & 1 & 0\\ 1 & a & b' & ab'\\ 1 & a' & b' & a'b \end{matrix}\right|=-(b-b')^2\left|\begin{matrix} 0 & 0 & 1 & 0\\ 0 & 0 & 1 & a\\ 1 & a & b' & ab'\\ 1 & a' & b' & a'b \end{matrix}\right|=$$ $$-(b-b')^2\left|\begin{matrix} 0 & 0 & a \\ 1 & a & ab'\\ 1 & a' & a'b \end{matrix}\right|=-(b-b')^2\left[a(a'-a)\right]=a(a-a')(b-b')^2 $$This is what I get. Please check If there is a mistake

OpenStudy (anonymous):

thanks for your help

OpenStudy (badhi):

you're welcome :)